I really need your help, Using Cramer's Rule, how would you solve this?
2x-y=4
3x+y-z=10
y+z=3

Answer by stanbon(75887)

Using Cramer's Rule, how would you solve this?
2x-y+0z=4
3x+y-z=10
0x+y+z=3
-------------------
The coefficient matrix is:
2..-1...0
3...1..-1
0...1...1
==================
The determinant of that matrix is D = 2+0+0-(0-2-3) = 7
---------------
The x-matrix is:
4..-1...0
10...1..-1
3...1...1
----
The determinant is Dx = 4+3-(0-4-10) = 21
---
Therefore x = Dx/D = 21/7 = 3
======================================
Form the y-matrix and find its determinant, Dy
Divide Dy by D to get "y".
------------------------------------
Form the z-matrix and find its determinant, Dz
Divide Dz by D to get "z".
=========================================
Cheers.
Stan H.